OpenStudy (anonymous):

what is the acceleration of a 1000 newton box on a 30 degrees frictionless ramp?

OpenStudy (shamim):

i think u know the equation\[F=mg\]

OpenStudy (shamim):

\[or, 1000=m \times10\]

OpenStudy (shamim):

or, m=100 kg

OpenStudy (shamim):

now we know that\[F=ma\]

OpenStudy (shamim):

here F will b thw force along the plane of ramp m= mass of the body a=acceleration

OpenStudy (shamim):

so\[1000 \times \cos 60=100a\]

OpenStudy (shamim):

if u need any more explanation plz ask me

OpenStudy (shamim):

just solve this equation for acceleration a

OpenStudy (ghazi):

this f=mg doesn't work here because its inclined plane

OpenStudy (anonymous):

funnily, though, since there is no friction, there is no need to take note of how heavy it is. a simply is a component of g  so, \(a=g sin 30\) or, \(a=10*0.5=5ms^{-2}\)